ADNAN MUHAMMED

Grade 12,

what will be the volume of a given mass of a gas at a pressure of 50 cm of hg if it occupies 260 mL at a pressure of 98 cm of Hg keeping the temperature constant

what will be the volume of a given mass of a gas at a pressure of 50 cm of hg if it occupies 260 mL at a pressure of 98 cm of Hg keeping the temperature constant
 

Grade:12th pass

2 Answers

Arun
6 years ago
Us
Arun
6 years ago
Uae boyle's law
P1 * V1 = P2 * V2
Put the values from the queations
98* 260 = 50 * V2
V2 = 98*260/50
V2 = 509.6 ml